# Extremum Seeking Tracking for Derivative-free Distributed Optimization

### v4

@@ V-B Monte Carlo simulations varying number of agents
Second, we test Extremum Seeking Tracking in a framework with $n=10$ and different number of agents, i.e., $N=5,10,20,30$. For each value of $N$, we generate communication graphs with a diameter $d$ such that the ratio $N/d$ is constant while varying $N$. Results are depicted in Fig. 3 show that the algorithm slows down as the number of agents increases.
Figure 3: Monte Carlo simulations for varying number of agents: mean and $1$-standard deviation band of cost error (left) and variable error (right).
@@ V-C Monte Carlo simulations varying the size $n$
We conclude this part by providing the results of a single trial in the case with $n=1$ to show the evolution of the solution estimate of each agent in error coordinate, with respect to the optimal solution to the problem.
@@ V-E Monte Carlo simulations for comparisons
In detail, Fig. 7 shows that, although the algorithm by [19] exhibits a faster convergence in the beginning phase of the simulations, our scheme has a better convergence rate and final accuracy. Indeed, although both Extremum Seeking Tracking and 1P-DSGT require a single function query for each iteration, we recall that 1P-DSGT adopts a stochastic dither whose theoretical gradient estimation properties are guaranteed only asymptotically and in the case of ergodic processes. On the opposite, our periodic dither guarantees that the theoretical gradient estimation is (practically) achieved with $\tau_{\text{per}}$ functions’ samples (in an averaged sense). This makes the convergence rate of our algorithm faster.
@@ VI Conclusions
The proof follows by setting $\ell_{i}(x_{i})=\dfrac{2}{\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}d_{i}^{k}R_{i,2}(x_{i},\delta d_{i}^{k})/\delta^{3}$. Finally, given a compact set $\mathcal{S}_{i}\subset{\mathbb{R}}^{n}$, let us bound $\left\|\ell_{i}(x_{i})\right\|$ for all $x_{i}\in\mathcal{S}_{i}$. Note that $\left\|\delta d_{i}^{t}\right\|\leq\delta\sqrt{n}$ for all $t\geq 0$ and let $\mathcal{S}^{\prime}_{i}\subset{\mathbb{R}}^{n}$ be a compact set such that (i) $\mathcal{S}_{i}\subseteq\mathcal{S}_{i}^{\prime}\subset{\mathbb{R}}^{n}$, and (ii) $x_{i}+\delta d_{i}^{t}\in\mathcal{S}^{\prime}$ for all $x_{i}\in\mathcal{S}_{i}$, $\delta\in(0,1]$, and $t\geq 0$. Thus, we can write
@@
where in $(a)$ we use the expression (46) of $R_{i,2}(x_{i},\delta d_{i}^{k})$, the definition of $\mathcal{S}_{i}^{\prime}$, and the fact that $\delta\in(0,1]$, while in $(b)$ we drop out the term $\delta^{3}$ from $(\delta d_{i}^{k})^{\alpha}$. We underline that, since the set $\mathcal{S}_{i}^{\prime}$ is compact and $f_{i}$ is smooth, $L_{i,\mathcal{S}_{i}}^{\prime}$ exists and is finite. The bound of $\ell_{i}(x_{i})$ follows by defining $L_{i,\mathcal{S}_{i}}:=2L_{i,\mathcal{S}_{i}}^{\prime}\sqrt{n}$ and combining the result (48) with the bound about the norm of the dither signal, i.e., $\left\|d_{i}^{t}\right\|\leq\sqrt{n}$ for all $t\geq 0$.
@@
In detail, [38, Th. 3.1] introduces $\bar{T}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ to define $\bar{\xi}:=\bar{T}\xi$, and a matrix $\bar{P}=\bar{P}^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ such that
@@
for all $\xi_{\text{avg}}\in{\mathbb{R}}^{n_{\xi}}$ and some $\bar{a}_{1},\bar{a}_{2},\bar{a}_{3}>0$. Let $P:=\bar{T}^{-1}\bar{P}T$. Then, there exist $a_{1},a_{2},a_{3}>0$ such that
@@
$\Delta V(\xi_{\text{avg}}^{t})<-\gamma c_{1}^{\prime}\left\|\xi_{\text{avg}}^{t}\right\|^{2}.$ (57)
Thus, the inequality (57) ensures that the set $\Omega_{r_{\xi}}$ is invariant for system (27). Hence, if we pick $\xi_{\text{avg}}^{0}\in\mathcal{B}_{r_{\xi}}$, we prove that $\xi_{\text{avg}}^{t}\in\Omega_{r_{\xi}}$ for all $t\geq 0$. Consequently, the bound (54) holds for all $t\geq 0$ and, in turn, also the inequality (57) is verified for all $t\geq 0$, namely we proved that the trajectories of system (27) enter the ball $\mathcal{B}_{\rho}$ exponentially fast. The result (30) follows from the inequality (57) and (29a) by setting $c_{3}:=c_{1}^{\prime}/(2a_{2})$.

### v6

@@ V-B Monte Carlo simulations varying number of agents
Second, we test Extremum Seeking Tracking in a framework with $n=10$ and different number of agents, i.e., $N=5,10,20,30$. For each value of $N$, we generate communication graphs with a diameter $d$ such that the ratio $N/d$ is constant while varying $N$. The results depicted in Fig. 3 show that the algorithm slows down as the number of agents increases.
Figure 3: Monte Carlo simulations for a varying number of agents: mean and $1$-standard deviation band of cost error (left) and variable error (right).
@@ V-C Monte Carlo simulations varying the size $n$
We conclude this part by providing in Fig. 5 the results of a single trial in the case with $n=1$ to show the evolution of the solution estimate of each agent in error coordinate with respect to the optimal solution to the problem.
@@ V-E Monte Carlo simulations for comparisons
In detail, Fig. 7 shows that, although the algorithm by [19] exhibits a faster convergence in the beginning phase of the simulations, our scheme has a better convergence rate and final accuracy. Both 1P-DSGT and Extremum Seeking Tracking employ a single function query per agent to approximate the gradient. These methods rely on the concept that averaging across iterations around the “quasi-static” local solution estimate $x_{i}^{t}$ provides an accurate gradient approximation. However, in the context of their theoretical frameworks, 1P-DSGT attains this approximation using an infinite number of samples (mean of an ergodic process), whereas Extremum Seeking Tracking only needs $\tau_{\text{per}}$ samples. This makes the convergence rate of our algorithm faster.
@@ VI Conclusions
The proof follows by setting $\ell_{i}(x_{i})=\dfrac{2}{\tau_{\text{per}}}\sum_{k=t+1}^{t+\tau_{\text{per}}}d_{i}^{k}R_{i,2}(x_{i},\delta d_{i}^{k})/\delta^{3}$. Finally, given a compact set $\mathcal{S}_{i}\subset{\mathbb{R}}^{n}$, let us bound $\left\|\ell_{i}(x_{i})\right\|$ for all $x_{i}\in\mathcal{S}_{i}$. Note that $\left\|\delta d_{i}^{t}\right\|\leq\delta\sqrt{n}$ for all $t\in\mathbb{N}$ and let $\mathcal{S}^{\prime}_{i}\subset{\mathbb{R}}^{n}$ be a compact set such that (i) $\mathcal{S}_{i}\subseteq\mathcal{S}_{i}^{\prime}\subset{\mathbb{R}}^{n}$, and (ii) $x_{i}+\delta d_{i}^{t}\in\mathcal{S}^{\prime}$ for all $x_{i}\in\mathcal{S}_{i}$, $\delta\in(0,1]$, and $t\in\mathbb{N}$. Thus, we can write
@@
where in $(a)$ we use the expression (46) of $R_{i,2}(x_{i},\delta d_{i}^{k})$, the definition of $\mathcal{S}_{i}^{\prime}$, and the fact that $\delta\in(0,1]$, while in $(b)$ we drop out the term $\delta^{3}$ from $(\delta d_{i}^{k})^{\alpha}$. We underline that, since the set $\mathcal{S}_{i}^{\prime}$ is compact and $f_{i}$ is smooth, $L_{i,\mathcal{S}_{i}}^{\prime}$ exists and is finite. The bound of $\ell_{i}(x_{i})$ follows by defining $L_{i,\mathcal{S}_{i}}:=2L_{i,\mathcal{S}_{i}}^{\prime}\sqrt{n}$ and combining the result (48) with the bound about the norm of the dither signal, i.e., $\left\|d_{i}^{t}\right\|\leq\sqrt{n}$ for all $t\in\mathbb{N}$.
@@
In detail, [38, Th. 3.1] introduces a full-rank matrix $\bar{T}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ to define $\bar{\xi}:=\bar{T}\xi$, and a matrix $\bar{P}=\bar{P}^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ such that
@@
for all $\xi_{\text{avg}}\in{\mathbb{R}}^{n_{\xi}}$ and some $\bar{a}_{1},\bar{a}_{2},\bar{a}_{3}>0$. Let $P:=\bar{T}^{\top}\bar{P}T$. Then, there exist $a_{1},a_{2},a_{3}>0$ such that
@@
$\Delta V(\xi_{\text{avg}}^{t})\leq-\gamma c_{1}^{\prime}\left\|\xi_{\text{avg}}^{t}\right\|^{2}.$ (57)
Thus, the inequality (57) ensures that the set $\Omega_{r_{\xi}}$ is forward-invariant for system (27). Hence, if we pick $\xi_{\text{avg}}^{0}\in\mathcal{B}_{r_{\xi}}$, we prove that $\xi_{\text{avg}}^{t}\in\Omega_{r_{\xi}}$ for all $t\in\mathbb{N}$. Consequently, the bound (54) holds for all $t\in\mathbb{N}$ and, in turn, also the inequality (57) is verified for all $t\in\mathbb{N}$, namely we proved that the trajectories of system (27) enter the ball $\mathcal{B}_{\rho}$ exponentially fast. The result (30) follows from the inequality (57) and (29a) by setting $c_{3}:=c_{1}^{\prime}/(2a_{2})$.